three numbers are chosen at random without replacement from {1,2,3,....10}.the probability that the minimum of the chosen numbers in 3 or their maximum is 7 is

Grade:12

1 Answers

Arun
6 years ago
 
Number of elements in sample space = 10C=120 
According to the question 3 & 7 must get chosen and the third no. should be between them (ie 4 or5 or 6)as per condition .
So, no of favourable cases =  3C1 =3
Therefore required probablity = 3/120 = 1/40